OpenStudy (anonymous):

Which functions have a positive rate of change and which have a negative rate of change? 2x – 3y = 9 A. positive B. negative

OpenStudy (susanna):

put it in y=mx+b form

OpenStudy (susanna):

2x-3y=9 move 2x to the other side - it becomes negative. -3y=-2x+9 divide both sides by -3 \[y =\frac{ 2 }{ 3 }x-3\]

OpenStudy (cal.lavender):

Do you know what rate of change is equal to?

OpenStudy (anonymous):

i think so. been a while since i actually did it doe

OpenStudy (susanna):

slope is positive, so the rate of change is positive

OpenStudy (cal.lavender):

It is equal to slope as she said

OpenStudy (anonymous):

ah. okay good

OpenStudy (anonymous):

y = −4x + 5 positive or negative

OpenStudy (susanna):

slope is negative, so....

OpenStudy (anonymous):

yeah was thinkig the same

OpenStudy (anonymous):

x y −1 −5 1 1 3 7 think this one is negative too

OpenStudy (susanna):

find the slope and see if it's negative or positive

OpenStudy (anonymous):

oh wait. it may be positive

OpenStudy (cal.lavender):

you also need a table for that

OpenStudy (anonymous):

...it is a table

OpenStudy (anonymous):

hold on. i'll draw it

OpenStudy (susanna):

\[\frac{ 1+5 }{ 1+1 }=\frac{ 6 }{ 2 }=3\]

OpenStudy (susanna):

slope is positive, so the rate of change is...

OpenStudy (anonymous):